# Multi-Modal Graph Convolutional Network with Sinusoidal Encoding for Robust Human Action Segmentation

### v1

@@ V-C Ablation Study
For the efficiency analysis, our model archives an optimal balance between computational efficiency and performance, operating at $131.0$ GFLOPs, significantly lower than using higher visual temporal resolution. For instance: processing motion and visual modalities at a resolution ratio of $30:2$ increases FLOPs to $220.0$ G, while uniform high-resolution processing ($30:30$) escalates costs to $2687.3$ GFLOPs, a 20.5×20.5$\times$20.5 × increase. This exponential scaling highlights the inefficiency of naively aligning modalities at high frequencies. While our framework introduces moderate computational overhead compared to models using pure motion sequences, it retains real-time performance (81.0 FPS on a single GPU) by strategically decoupling temporal resolutions.
V-D Comparison with states-of-the-art
@@
The proposed MMGCN is compared with state-of-the-art methods in the field of human action segmentation on the Bimanual Actions dataset [24]. The compared methods include the method proposed by Dreher et al. [24], H2O+RGCN [25], BiRNN [26], 2G-GCN [27], PGCN [12] and TFGCN [1].
@@
The cross-validation results are listed in Table II. Our proposed MMGCN achieves state-of-the-art performance across all evaluation metrics, as validated on the leave-one-subject-out benchmark. Compared to TFGCN [1], the previous best method, MMGCN improves accuracy from $88.4\%$ to $89.2\%$ and F1 macro from $88.6\%$ to $89.1\%$, with significant gains in boundary-sensitive metrics: F1@10 increases from $93.7$ to $94.3$, F1@25 from $91.9$ to $92.7$, and F1@50 from $85.4$ to $85.9$. Compared to PGCN [12], which achieved $88.5\%$, $85.5\%$, and $77.0\%$ action segmentation performance, MMGCN demonstrates substantial improvements, particularly at higher overlap thresholds, indicating better temporal localization of action boundaries. These improvements highlight MMGCN’s robustness to temporal misalignment and the efficacy of the multi-modal fusion strategy. Despite achieving superior performance, MMGCN exhibits a slightly higher standard deviation compared to TFGCN [1], particularly at the F1@10 and F1@25. This increased variability might be attributed to differences in model architecture and sensitivity to variations in motion and image inputs. This observation motivates further experiments under noise-induced conditions to evaluate the model’s robustness, ensuring consistent performance across diverse scenarios.
@@
To evaluate the robustness of MMGCN under noisy conditions, we selected the most common noise - misdetection. Specifically, on the testset of subject one of the Bimanual Actions dataset [24], the motion nodes are randomly removed per frame. The removal rate (noise intensity) increases from $0\%$ to $25\%$. The results of robustness are illustrated in Fig. 4. As noise intensity increases, the gap in action (a) recognition and (b) segmentation accuracy between the MMGCN and the other two models progressively widens. Notably, the MMGCN consistently maintains the highest performance across all noise levels and, in the segmentation case, overtakes competing models to secure the top position. This trend highlights the superior resilience of our approach to partially missing inputs.
@@
Figure 3 presents the qualitative comparison of action segmentation results for the breakfast cereal preparation scenario from the Bimanual Actions dataset [24]. Among the compared models, our proposed MMGCN produces predictions that most closely align with the ground truth. It successfully mitigates the over-segmentation issues, which are commonly observed in other methods. This demonstrates the model’s effectiveness in maintaining coherent action segments. However, MMGCN overlooks actions with short durations or extends their intervals, leading to temporal shifts in action boundaries.
@@ ACKNOWLEDGMENT
We extend our sincere gratitude to Timothy Ju Kin, Ng for his invaluable technical insights and support.

### v2

@@ V-D Comparison with states-of-the-art
The proposed MMGCN is compared with state-of-the-art methods in the field of human action segmentation on the Bimanual Actions dataset [dreher2019learning]. The compared methods include the method proposed by Dreher et al. [dreher2019learning], H2O+RGCN [lagamtzis2023exploiting], BiRNN [morais2021learning], 2G-GCN [qiao2022geometric], PGCN [xing2022understanding] and TFGCN [xing2024understanding].
@@
The cross-validation results are listed in Table II. Our proposed MMGCN achieves state-of-the-art performance across all evaluation metrics, as validated on the leave-one-subject-out benchmark. Compared to TFGCN [xing2024understanding], the previous best method, MMGCN improves accuracy from $88.4\%$ to $89.2\%$ and F1 macro from $88.6\%$ to $89.1\%$, with significant gains in boundary-sensitive metrics: F1@10 increases from $93.7$ to $94.3$, F1@25 from $91.9$ to $92.7$, and F1@50 from $85.4$ to $85.9$. Compared to PGCN [xing2022understanding], which achieved $88.5\%$, $85.5\%$, and $77.0\%$ action segmentation performance, MMGCN demonstrates substantial improvements, particularly at higher overlap thresholds, indicating better temporal localization of action boundaries. These improvements highlight MMGCN’s robustness to temporal misalignment and the efficacy of the multi-modal fusion strategy. Despite achieving superior performance, MMGCN exhibits a slightly higher standard deviation compared to TFGCN [xing2024understanding], particularly at the F1@10 and F1@25. This increased variability might be attributed to differences in model architecture and sensitivity to variations in motion and image inputs. This observation motivates further experiments under noise-induced conditions to evaluate the model’s robustness, ensuring consistent performance across diverse scenarios.
@@
The observed average improvement of $1\%$ in Top-1 accuracy compared to state-of-the-art methods motivates further analysis of our approach’s robustness in real-world scenarios. To simulate realistic conditions, we progressively introduce input noise (simulating detection failures) by randomly removing motion nodes from both human joints and object centroids at each frame in Subject 1’s test set from the Bimanual Actions dataset [dreher2019learning]. We systematically increase the noise intensity (node removal rate) from $0\%$ to $25\%$. For each frame, we randomly select $V^{\prime}=\eta\times V$ nodes without replacement (where $\eta$ is the noise intensity ranging 0-$25\%$), setting all feature channels of selected nodes to zero. The results of robustness are illustrated in Fig. 5. As noise intensity increases, the gap in action (a) recognition and (b) segmentation accuracy between the MMGCN and the other two models progressively widens. Notably, the MMGCN consistently maintains the highest performance across all noise levels and, in the segmentation case, overtakes competing models to secure the top position. This trend highlights the superior resilience of our approach to partially missing inputs.
@@
Figure 4 presents the qualitative comparison of action segmentation results for the breakfast cereal preparation scenario from the Bimanual Actions dataset [dreher2019learning] with input frames illustrating critical perception artifacts. Frame (a) show correct detections, frames (b-c) exhibit object misdetections (missing bowl, cutting board), and frame (d) contains depth errors from hand-object occlusion (miscalculated bowl position). While comparative models (TFGCN, PGCN, ASSIGN) exhibit severe over-segmentation due to these input artifacts, particularly misaligning ”hold”, ”lift” and ”approach”, our MMGCN successfully mitigates the over-segmentation issues and maintains coherent segments closest to ground truth. This demonstrates the model’s effectiveness in maintaining coherent action segments. However, MMGCN overlooks actions with short durations or extends their intervals, leading to temporal shifts in action boundaries.
@@ ACKNOWLEDGMENT
The research has been (partially) supported by the German Research Foundation (DFG) SFB 1320 EASE, CRC, University of Bremen. The research was conducted in subproject R1: NEEM-based embodied knowledge system. We extend our sincere gratitude to Timothy Ju Kin, Ng for his invaluable technical insights and support.